# Multi-scale phylodynamic modelling of rapid punctuated pathogen evolution

### v2

@@ Abstract
Computational multi-scale pandemic modelling remains a major and timely challenge. Here we identify specific requirements for a new class of pandemic models operating across three scales: (1) rapid pathogen evolution, punctuated by emergence of new variants, (2) human interactions within a heterogeneous population, and (3) public health responses which constrain individual actions to control the disease transmission. We then present a pandemic modelling framework satisfying these requirements and capable of simulating multi-scale dynamic feedback loops. The developed framework comprises a stochastic agent-based model of pandemic spread, coupled with a phylodynamic model of the within-host pathogen evolution. It is validated with a case study, modelling a rapid punctuated evolution of SARS-CoV-2, based on global and contemporary genomic surveillance data, during the COVID-19 transmission within a large heterogeneous population. We demonstrate that the model captures the essential features of the COVID-19 pandemic and the novel coronavirus evolution, while retaining computational tractability and scalability.
@@ 1 Main
Digital epidemiology is an emerging field, rapidly capitalising on the increasing availability of high-resolution genomic, immuno-epidemiological, demographic and human mobility data, social media analytics, high-performance computing power, as well as innovations in simulation methods and data science. In principle, these diverse sources of data should allow modellers to develop informative multi-scale pandemic models with substantially higher accuracy and predictability. Yet, pandemic modelling continues to face significant challenges due to (i) the inherent complexity of rapidly evolving pathogens, (ii) population heterogeneity (demographic, immunological and behavioural), (iii) multi-objective public health interventions carried out under severe pressure and non-trivial social dynamics.
@@
The modelling complexity is exacerbated by (iv) different time scales needed to model pandemics on multiple levels, ranging from within-host changes in implicated pathogens, to natural infections in individuals and their social behaviour and interactions (v) fragmentation of data across heterogeneous sources, and (vi) computational complexity of multiple simulations over a sufficiently long horizon, required to establish statistically significant outcomes. Multi-scale models often suffer from the “curse of dimensionality”, when computational costs increase exponentially with the number of degrees of freedom [1].
@@
A principal modelling problem is the presence of feedback loops: for example, pandemic mitigation measures may put an evolutionary pressure on the pathogen, causing the emergence of more transmissible lineages. Higher transmissibility may increase the need for more vigorous interventions, which in turn may cause changes in how populations respond and behave, and increase evolutionary pressure on the pathogen. One possible outcome of this feedback loop is the establishment of recurrent waves and a nonlinear transition from emerging to endemic pathogen. Consequently, multi-scale modelling of a major pandemic crisis, such as the COVID-19, quickly becomes intractable.
@@
Over the last decades, stochastic agent-based modelling (ABM) has been established as a robust tool for tracing fine-grained effects of complex intervention policies in diverse epidemic and pandemic settings [2, 3, 4, 5]. Most recently, these studies produced policy recommendations developed for COVID-19 control, and adopted in Australia [6, 7, 8], the USA [9], the UK [10], and broadly by the WHO [11]. A largely unexplored avenue to leverage the precision and fidelity of ABMs is to extend them with comprehensive phylodynamic modelling of evolving pathogens, going beyond existing phylogenetic models which define simplified evolutionary landscapes [12, 13]. This necessitates a new class of multi-scale phylodynamic ABMs.
@@
An optimal framework for multi-scale phylodynamic agent-based modelling should include several distinct capabilities that produce quantifiable outcomes. These include the ability to:
  1. 1.
@@
      Trace changes in transmissibility with respect to pathogen mutations. This includes analysis of the average basic reproductive ratio ($R_{0}$) within the population, and the accumulated mutations measured by the average genomic distance between a circulating genome and the ancestral genome, i.e., in terms of the substitution rate (the number of non-synonymous mutations per year observed in detected cases).
@@
      Examine the genomic structure-function dynamics. This analysis quantifies the individual contributions of amino acids to changes in the pathogen’s fitness, and traces the dynamics of these contributions over time.
  3. 3.
@@
      Explore temporal alignments between phylodynamics and disease dynamics, e.g., relating accumulated mutations, genomic diversity, and saltations in pathogen fitness to incidence peaks.
@@
A robust phylodynamic ABM needs to be validated by comparing its target outcomes against the ground truth dynamics. Once validated, it can be used to explore diverse counter-factual scenarios with respect to phylodynamic, demographic and immuno-epidemiological characteristics. We illustrate the ground truth dynamics, matching the three capabilities, in Figures 1 to 4, by using available genomic and disease surveillance data on COVID-19 and SARS-CoV-2 from 2020 to 2024.
To demonstrate Capability 1, a modelling framework is expected to generate pandemic or epidemic patterns aligned with ground truth (i.e., observed data). For example, Fig. 1 shows patterns observed in global data reported during the COVID-19 pandemic, highlighting prominent peaks and recurrent incidence waves. We point out that each incidence peak is temporally aligned with the emergence of a new variant of concern. Notably, the two most prominent incidence peaks occurred in early 2022 and early 2023, corresponding to the dominance of Omicron BA.1 and Omicron XBB, respectively. These observations suggest a complex interplay between pandemic patterns and evolutionary phylodynamic features of the viral variants in circulation.
@@
Capability 2 demands a dynamic analysis of the pathogen fitness in terms of observable phylodynamic characteristics. For example, Fig. 2 relates the growing transmissibility of SARS-CoV-2 to (i) the mutations accumulated relative to its ancestral strain, and (ii) the associated genomic diversity of the evolved pathogen. It also places the fitness dynamics in the context of the genomics.
@@
Specifically, Fig. 2 (A) shows a rapid punctuated increase in fitness, with two significant surges in the relative transmissibility (measured relative to the basic reproductive ratio of the ancestral strain $R_{A}$), and the accumulated mutations, observed during early 2021 and early 2022. The accumulated mutations continued to grow after 2022, reaching 130 substitutions by mid-2024 at the rate of roughly 30 substitutions per year, according to linear regression (Fig. 2 (B)).
@@
Furthermore, notable changes in the amino acid presence were observed across the genome, particularly the spike region, between the start of the pandemic and late 2023, as illustrated for the 5 end of the S gene in Fig. 2 (C–D). These shifts occurred preferentially at specific positions, deviating significantly from the ancestral genome, and thus, potentially contributing to the increase in fitness.
@@
Figure 2: Capability 2. Pathogen fitness with respect to its phylodynamics. (A) Punctuated growth of fitness, measured relative to the basic reproductive ratio of the ancestral strain $R_{A}$ [16] (represented as the $R/R_{A}$, red circles), overlaid with the average accumulated mutations across the population (represented by the Hamming distance $\widehat{D}$), also measured relative to the ancestral strain. (B) Accumulated mutations $\widehat{D}$, expressed as the number of substitutions relative to the ancestral genome. The black line shows the linear regression of the accumulated mutations ($\widehat{D}$) between 2020 and 2024, with a gradient of roughly 30 substitutions per year. We refer to Fig. 1 for clade colours. For visualisation purposes, we show around 4,000 genomes [17]. (C, D) Histograms of amino acids over the first 26 codons in gene S of the COVID-19 ancestral strain (C), and 631 randomly selected genomes from GenBank and the Robert Koch Institute between 17 and 31 December 2023 (D) [15].
@@
Capability 3 is focused on the emergence and dominance of variants of concern, in the context of phylodynamic and epidemiological dynamics. For example, Fig. 3 (A) traces a continuous increase in the accumulated mutations which is contrasted with the fluctuating genomic diversity. Unlike the mutations $\widehat{D}$ which accumulated relative to the ancestral strain, the genomic diversity, measured as the pairwise distance $\overline{D}$ among the genomes of co-circulating pathogens, shows only a marginal increase during the observed period. We extracted phylodynamic features using the data recorded until early 2024, noting a significant drop in the number of sequenced samples worldwide from 2024 onwards (Fig. 3 (B)).
@@
Figure 3: Capability 3 (i). Temporal alignments between phylodynamics and disease dynamics. (A) The global COVID-19 incidence (cases per million, shaded blue area) overlaid with accumulated mutations $\widehat{D}$ (solid brown line) and genomic diversity $\overline{D}$ (represented by the average pairwise distance between two randomly selected genome sequences, dashed brown line). (B) The number of sequences, measured using a 7-day moving average, captured by the open SARS-CoV-2 sequence database processed by Nextstrain [15] (shaded orange area) overlaid with accumulated mutations $\widehat{D}$ across the genome (solid brown line), and genomic diversity $\overline{D}$ (dashed brown line).
@@
In general, abrupt changes in genomic diversity are related to the frequency of different variants reported during the pandemic period (Fig. 4 (A)), and considered in context of the corresponding phylogenetic tree (Fig. 4 (B)). For example, we observed that during the rapid evolution of SARS-CoV-2, sudden decreases in circulating diversity correspond to specific lineages becoming dominant, whereas new variants are more likely to emerge during periods of increasing circulating diversity (i.e., increasing pairwise genomic distance). Importantly, the observed changes in the pairwise distance are also reflected in the phylogenetic tree produced from representative sub-sampling of global sequences (Fig. 4 (B)). Branches corresponding to new variants of concern tend to descend from more basal lineages than the main lineages circulating immediately prior to their emergence. These evolutionary saltations may explain step changes in transmissibility and virulence.
@@
Figure 4: Capability 3 (ii). Emergence and dominance of variants of concern. (A) Genomic diversity (solid black line) overlaid with the clade frequency between 2020 and 2024. (B) Phylogenetic tree of SARS-CoV-2, built with Nextstrain from worldwide sequences collected between 2020 and 2024 [17]. The phylogenetic tree is constructed from approximately 4,000 genomes.
@@ 2 Results
In this section, we overview the model’s multi-scale approach (subsection 2.1). We then apply PhASE TraCE to a case study of SARS-CoV-2, and evaluate the simulated phylodynamics against the modelling capabilities and objectives (subsection 2.2). Finally, we explore counterfactual scenarios by varying specific assumptions of the case study, e.g., the role of chronic infections and population sizes (Supplementary section F).
@@ 2.1 Model overview
This multi-scale feedback is realised by incorporating four concurrent dynamics implemented in distinct processing layers: demographic, epidemic, immunological, and phylogenetic, as illustrated in Fig. 5. The demographic layer defines the population structure (i.e., host type) and social groups constraining agent interactions. The transmission is simulated by the epidemic layer, tracing individual interactions occurring within and across different social contexts, including residential (e.g., household) and professional settings (e.g., working group). In addition, the epidemic layer sets out an intervention scenario, with varying NPI adoption levels and vaccination strategies, thus reflecting changes in health policy and public opinion [26, 27]. The infection transmission depends on the agents’ immunity levels derived from their individual dynamic history of exposures and immunisation, determined by the immunological layer. There are two infected host categories: typical infected hosts and chronically infected hosts, with the latter representing hosts with persistent SARS-CoV-2 infections arising due to some underlying factors and resulting in longer recovery periods [52, 46] (see Supplementary section A.4). The phylogenetic layer determines the within-host evolution of pathogens in terms of mutation and selective pressure. Section 4 provides a detailed description and implementation of these four layers.
Figure 5: Model overview. The demographic layer assigns agents to social contexts, including household, household cluster, neighbourhood, community, and working group/school. The immunological layer keeps track of the agents’ immunisation records, infection histories, and the host types. The epidemic layer describes an intervention scenario, in terms of NPIs and their adoption levels: case isolation (CI), home quarantine (HQ), social distancing (SD), and school closure (SC), along with a dynamic immunisation schedule. Both epidemic and phylogenetic layers affect inter-host transmission (e.g., on day $n$, a chronically infected host $j$ infects their family member $i$ and passes genome $s_{j}(n)$ to agent $i$). The phylogenetic layer determines the genomic mutations under various strengths of selective pressure, and traces changes in transmissibility (reflected in $R_{0}$). The fitness of each genome is determined by a weight table, quantifying the contributions of each amino acid in a given codon position.
@@ 2.2 Case study: Rapid punctuated evolution of SARS-CoV-2
PhASE TraCE was calibrated to the COVID-19 pandemic and the evolutionary trajectory of SARS-CoV-2 over the period of four years, from 2020 to 2023. The results comprise simulations over a six-year period (2020 — 2025), across three artificial agent populations: 1.7 million agents (small), 8 million agents (medium), and 25.4 million agents (large), as described in Supplementary Section B.
In calibrating PhASE TraCE, we examined COVID-19 pandemic and the size of susceptible, recovered and vaccinated populations (Capability 1); phylodynamic characteristics including transmissibility, fitness, the number of accumulated mutations relative to the ancestral strain, and genomic diversity (Capability 2); and the emergence of variants, analysed using the phylogenetic tree and statistical methods (Capability 3). The robustness of the model is established through sensitivity analysis, by varying key parameters and assessing their impact on simulation outcomes. See Supplementary Table 4 for the list of phylogenetic parameters used in the SARS-CoV-2 case study and Section G for more information on sensitivity analysis.
@@
Capability 1. Our results show recurrent incidence waves as illustrated in Fig. 6 (A), in concordance with empirical observations (Fig. 1). Although the first simulated incidence peak (detected during 2021) is noticeably higher in larger populations, the incidence converged to around 2,000 cases per million after 2023, indicating endemicity in all simulated populations.
The main contributing factor to the recurrent incidence is the fluctuating number of recovered individuals (Fig. 6 (B)) that replenish the susceptible population after their immunity wanes. This subsequently leads to an increasing number of re-infection cases. Additionally, due to the diminishing immunity, declining vaccine uptake, and declining adoption of social distancing in the population (simulated to be decreasing from 50% of hosts in 2020 to 20% from 2024 onwards), the overall incidence slightly increases over the simulated timeframe after the first peak (Fig. 6 (A)).
We also observe an increasing number of individuals with multiple vaccination and infection records (Fig. 6 (C)). These dynamics slow down the transmission but are unable to completely eliminate the spread, due to the waning immunity and reduced vaccine efficacy against mutated variants. Figures 6 (B) and (C) only show dynamics for the 1.7 million population. We note that the pandemic patterns are consistent across different population sizes, as shown in Supplementary Fig. 19.
@@
Importantly, we observe a concurrent punctuated increase (i.e., ‘jump’) in both transmissibility $R_{0}$, i.e., fitness $K$ (Fig. 7 (A)), and accumulated mutations $\widehat{D}$ (Fig. 7 (B)) at the start of 2021. This is a pattern which is also observed in empirical data illustrated in Fig. 2 (A). The timing of the simulated ‘jump’ aligns well with the first incidence peak in 2021, illustrated in Fig. 6 (A). This alignment can be explained by the accumulation of fitness-increasing mutations in chronically infected hosts, due to a higher selective pressure on pathogens evolving in these hosts; consequently generating a highly transmissible strain (i.e., a variant of concern) (see Supplementary Section A.4). This is confirmed by a comparison with a counterfactual scenario without chronically infected hosts in which the jump is not observed, as shown in Fig. 26 and analysed in Supplementary Section F.
Similar jumps are observed at later pandemic stages in a majority of individual simulation runs (see Fig. 7 (C) and Supplementary Figures 20 to 22), but their timing differs across realisations, and so the average profiles shown in Fig. 7 (A) and (B) smooth the saltations. The very first saltation is mostly aligned across the individual realisations, and so it is quite prominent in both $K$ and $\widehat{D}$. Figure 7 (C) shows that the fitness $K$ and accumulated mutations $\widehat{D}$ are strongly temporally aligned (with Pearson correlation $0.99$ over $2,191$ data points), suggesting that saltations in transmissibility are produced by accumulated mutations.
@@
To explain the punctuated increase in fitness, we traced changes in the distribution of fitness contributions across the simulated genome (Capability 2(ii)) by comparing the relative frequencies of amino acids in simulated ancestral genome (Fig. 8 (I)) and the evolved distributions in circulating genomes after six simulation years (Fig. 8 (II)). In simulated dynamics, we observed a clear shift towards amino acids with positive fitness contributions, in accordance with empirical observations (Fig. 2 (E) and (F)). This observation indicates that the viral mutations are subject to selective pressure transitioning to a higher point in the viral fitness landscape. Section 4 and Supplementary section A.3 provide a detailed description of the fitness contribution method.

### v3

@@ Abstract
Computational multi-scale pandemic modelling remains a major and timely challenge. Here we identify specific requirements for a new class of models simulating pandemics across three scales: (1) pathogen evolution, often punctuated by the rapid emergence of new variants, (2) human interactions within a heterogeneous population, and (3) public health responses which constrain individual actions to control the disease transmission. We then present a pandemic modelling framework satisfying these requirements and capable of simulating feedback loops between dynamics unfolding at these different scales. The developed framework comprises a stochastic agent-based model of pandemic spread, coupled with a phylodynamic model that incorporates within-host pathogen evolution. It is validated with a case study, modelling the punctuated evolution of SARS-CoV-2, based on global and contemporary genomic surveillance data, which captures a large heterogeneous population. We demonstrate that the model replicates the essential features of the COVID-19 pandemic and virus evolution, while retaining computational tractability and scalability.
@@ 1 Introduction
Digital epidemiology is an emerging field, rapidly capitalising on the increasing availability of high-resolution genomic, immuno-epidemiological, demographic and human mobility data, social media analytics, high-performance computing power, as well as innovations in simulation methods and data science. In principle, these diverse sources of data should allow modellers to develop informative multi-scale pandemic models with a better capacity (i) to simulate realistic epidemiological, immunological and evolutionary dynamics, and (ii) to anticipate longer-term epidemiological and evolutionary dynamics as the pandemic unfolds. Yet, pandemic modelling continues to face significant challenges in concisely capturing relevant characteristics of pandemic pathogens, such as their pathogenesis, transmissibility, antigenicity, etc., as well as predicting long-term phylodynamic trajectories and future public health risks. These challenges arise due to (i) the inherent complexity of rapidly evolving pathogens, (ii) population heterogeneity (demographic, immunological and behavioural), and (iii) multi-objective public health interventions carried out under severe pressure and non-trivial social dynamics.
@@
The modelling complexity is exacerbated by (iv) different time scales needed to model pandemics on multiple levels, ranging from the range of evolutionary drivers and rates of the implicated pathogens to natural infections in individuals and their social behaviour and interactions, (v) fragmentation of data across heterogeneous sources, and (vi) computational complexity of multiple simulations over a sufficiently long horizon, required to examine the distribution of outcomes in many stochastic realisations of the model and across ranges of uncertainties in parameters such as substitution rate, fitness, accumulated mutations, and genomic diversity. Multi-scale models often suffer from the “curse of dimensionality”, when computational costs increase exponentially with the number of degrees of freedom [1].
@@
A principal modelling problem is the presence of feedback loops: for example, pandemic mitigation measures may indirectly affect the pathogen evolution, leading to the emergence of more transmissible lineages. Higher transmissibility may increase the need for more vigorous interventions, which in turn may cause changes in how populations respond and behave, and constrain the pathogen evolution in a specific way. This feedback loop contributes to the formation of recurrent waves of infection, fluctuating genomic diversity, non-linear increases in fitness levels, and a potentially delayed transition to endemicity. Consequently, multi-scale modelling of a major pandemic crisis, such as COVID-19, quickly becomes intractable.
@@
Over the last decades, stochastic agent-based modelling (ABM) has been established as a robust tool for tracing fine-grained effects of complex intervention policies in diverse epidemic and pandemic settings [2, 3, 4, 5]. Most recently, these studies produced policy recommendations developed for COVID-19 control, which were adopted in Australia [6, 7, 8], the USA [9], the UK [10], and broadly by the WHO [11]. In these ABMs, each agent represents an individual human host with a set of demographic, epidemiological, and immunological attributes.
A largely unexplored avenue to leverage the precision and fidelity of ABMs is to extend them with comprehensive phylodynamic modelling of evolving pathogens, going beyond existing phylogenetic models which define simplified evolutionary landscapes [12, 13]. This necessitates a new class of multi-scale phylodynamic ABMs.
@@
An effective framework for multi-scale phylodynamic agent-based modelling should include the following distinct capabilities that produce quantifiable outcomes:
  1. Capability 1:
@@ 2. Capability 2:
      Trace changes in transmissibility with respect to pathogen mutations. This includes analysis of the average reproductive number within the population, and the accumulated mutations measured by the average genomic distance between a circulating genome and the ancestral genome, e.g., in terms of the nucleotide substitution rate.
@@
      Examine the functional dynamics (i.e., the relationship between the genome sequence and the pathogen’s epidemic behaviour). This analysis quantifies the individual contributions of amino acids to changes in the pathogen’s fitness, and traces the dynamics of these contributions to transmissibility and antigenicity over time.
  3. Capability 3:
@@
      Explore concordance between phylodynamics and disease dynamics, e.g., relating accumulated mutations, genomic diversity, and saltations in pathogen fitness to incidence peaks.
@@
A robust phylodynamic ABM needs to be validated by comparing its target outcomes against the ground truth dynamics. Once validated, it can be used to explore diverse counterfactual scenarios with respect to phylodynamic, demographic and immuno-epidemiological characteristics. We illustrate the ground truth dynamics, matching the three capabilities, in Figs 1 to 4, by using available genomic and disease surveillance data on SARS-CoV-2 and COVID-19 respectively from 2020 to 2024 (we note that the detected incidence is affected by the testing capacity).
To demonstrate Capability 1, a modelling framework is expected to generate pandemic or epidemic patterns aligned with ground truth (i.e., observed data). For example, Fig 1 shows patterns observed in global data reported during the COVID-19 pandemic, highlighting prominent peaks and recurrent incidence waves. We point out that each incidence peak is temporally aligned with the emergence of a new variant of concern. Notably, the two most prominent incidence peaks occurred in early 2022 and early 2023, corresponding to the dominance of Omicron BA.1 and Omicron XBB, respectively. These observations suggest a complex interplay between pandemic patterns and evolutionary phylodynamic features of the viral variants in circulation.
@@
Capability 2 demands a dynamic analysis of the pathogen fitness in terms of observable phylodynamic characteristics. For example, Fig 2 relates the growing transmissibility of SARS-CoV-2 to (i) the mutations accumulated relative to its ancestral strain, and (ii) the associated genomic diversity of the evolved pathogen. It also places the fitness dynamics in the context of the genomics.
@@
Punctuated evolution of the novel coronavirus was observed even during the first year of the pandemic [16]. Fig 2A shows a rapid punctuated increase in fitness for the first two years of the pandemic, with two significant surges in the relative transmissibility (measured relative to the basic reproductive number of the ancestral strain $R_{A}$), and the accumulated mutations, observed during early 2021 and early 2022. The accumulated mutations continued to grow after 2022, reaching 130 substitutions by mid-2024 at the rate of roughly 30 substitutions per year, according to linear regression (Fig 2B).
@@
Furthermore, notable changes in the amino acid presence were observed across the genome, particularly the spike region, between the start of the pandemic and late 2023, as illustrated for the 5 end of the S gene in Fig 2C–D. These shifts occurred preferentially at specific positions, deviating significantly from the ancestral genome, and thus, potentially contributing to the increase in fitness.
@@
Figure 2: Capability 2. Pathogen fitness with respect to its phylodynamics. (A) Punctuated growth of fitness, measured relative to the basic reproductive number of the ancestral strain $R_{A}$ [17] (represented as the $R/R_{A}$, red circles), overlaid with the average accumulated nucleotide substitutions across the population (represented by the Hamming distance $\widehat{D}$), also measured relative to the ancestral strain. (B) A time-scaled phylogeny of representative SARS-CoV-2 sequences mapped onto the number of accumulated substitutions relative to the ancestral genome. We refer to Fig 1 for clade colours. For visualisation purposes, we show around 4,000 genomes. Data retrieved from Nextstrain [18]. (C, D) Heatmaps of amino acids over the first 25 codons in gene S of the COVID-19 ancestral strain (C), and 631 randomly selected genomes from GenBank and the Robert Koch Institute between 17 and 31 December 2023 (D) [15].
@@
Capability 3 is focused on the emergence and dominance of variants of concern, in the context of phylodynamic and epidemiological dynamics. Two quantities can be observed on any given day: the average mutations accumulated by the evolved genomes relative to the ancestral strain, denoted $\widehat{D}$; and the average pairwise distance among the genomes of co-circulating pathogens on that day, denoted $\overline{D}$ (see Section 4). For example, Fig 3A traces a continuous increase in the accumulated mutations which is contrasted with the fluctuating genomic diversity. Unlike the mutations $\widehat{D}$ which accumulated relative to the ancestral strain, the genomic diversity, measured as the daily pairwise distance $\overline{D}$, shows only a marginal increase during the observed period. We extracted phylodynamic features using the data recorded until early 2024, noting a significant drop in the number of sequenced samples worldwide from 2024 onwards (Fig 3B).
@@
Figure 3: Capability 3 (i). Temporal alignments between phylodynamics and disease dynamics. (A) The global COVID-19 incidence (new weekly cases per million, shaded blue area) overlaid with accumulated mutations $\widehat{D}$ (represented by the average distance between circulating genomes and the ancestral strain, solid brown line) and genomic diversity $\overline{D}$ (represented by the average pairwise distance between two randomly selected genome sequences, dashed brown line). (B) The number of sequences, measured using a 7-day moving average, captured by the open SARS-CoV-2 sequence database processed by Nextstrain [15] (shaded orange area) overlaid with accumulated mutations $\widehat{D}$ across the genome (solid brown line), and genomic diversity $\overline{D}$ (dashed brown line).
@@
In general, abrupt changes in genomic diversity are related to the frequency of different variants reported during the pandemic period (Fig 4A), and considered in context of the corresponding phylogenetic tree (Fig 4B). For example, we observed that during the rapid evolution of SARS-CoV-2, sudden decreases in circulating diversity correspond to specific lineages becoming dominant, whereas new variants are more likely to emerge during periods of increasing circulating diversity (i.e., increasing pairwise genomic distance). Importantly, the observed changes in the pairwise distance are also reflected in the phylogenetic tree produced from representative sub-sampling of global sequences (Fig 4B). Branches corresponding to new variants of concern tend to descend from more basal lineages than the main lineages circulating immediately prior to their emergence. These evolutionary saltations may explain step changes in transmissibility and virulence.
@@
Figure 4: Capability 3 (ii). Emergence and dominance of variants of concern. (A) Genomic diversity $\overline{D}$ (i.e., the average pairwise distance between two randomly selected genome sequences on a given day, solid black line) overlaid with the clade frequency between 2020 and 2024. (B) Phylogenetic tree of SARS-CoV-2, built with Nextstrain from worldwide sequences collected between 2020 and 2024 [18]. The phylogenetic tree is constructed from approximately 4,000 genomes.
@@ 2 Results
In this section, we overview the model’s multi-scale approach (subsection 2.1). We then apply PhASE TraCE to a case study of SARS-CoV-2, and evaluate the simulated phylodynamics against the modelling capabilities and objectives (subsection 2.2). Finally, we explore counterfactual scenarios by varying specific assumptions of the case study, e.g., the role of chronic infections and population sizes (S1 Text: Counterfactual modelling).
@@ 2.1 Model overview
This multi-scale feedback is realised by incorporating four concurrent dynamics implemented in distinct processing layers: demographic, epidemic, immunological, and phylogenetic, as illustrated in Fig 5. The demographic layer defines the population structure (i.e., host type) and social groups constraining agent interactions. The transmission is simulated by the epidemic layer, tracing individual interactions occurring within and across different social contexts, including residential (e.g., household) and professional settings (e.g., working group). In addition, the epidemic layer sets out an intervention scenario, with varying NPI adoption levels and vaccination strategies, thus reflecting changes in health policy and public opinion [27, 28]. The infection transmission depends on the agents’ immunity levels derived from their individual dynamic history of exposures and immunisation, determined by the immunological layer. There are two infected host categories: typical infected hosts and chronically infected hosts, with the latter representing hosts with persistent SARS-CoV-2 infections arising due to some underlying factors and resulting in longer recovery periods [29, 30] (see S1 Text: Infected host categories). The phylogenetic layer determines the within-host evolution of pathogens in terms of mutation and selective pressure. Section 4 provides a detailed description and implementation of these four layers.
Figure 5: The demographic layer assigns agents to social contexts, including household, household cluster, neighbourhood, community, and working group/school. The immunological layer keeps track of the agents’ immunisation records, infection histories, and the host types. An example of the infectivity profiles is illustrated for a typical infected host (red line) and a chronically infected host (purple line). On day $n$, the infectivity peaks (i.e., reaches 1.0) for both host types (black dot). On day $n^{\prime}>n$, the infectivity of the chronically infected host (purple dot) reduces more slowly than that of the typical infected host (red dot). Upon recovery, the typical infected host attains perfect immunity of 1.0 for 60 days, after which the immunity starts to decline (green line). The epidemic layer describes an intervention scenario, in terms of NPIs and their adoption levels: case isolation (CI), home quarantine (HQ), social distancing (SD), and school closure (SC), along with a dynamic immunisation schedule. Both epidemic and phylogenetic layers affect inter-host transmission (e.g., on day $n$, a chronically infected host $j$ infects their family member $i$ and passes genome $s_{j}(n)$ to agent $i$). The phylogenetic layer determines the genomic mutations under various strengths of selective pressure, and traces changes in transmissibility (reflected in reproductive number). The fitness of each genome is determined by a weight table, quantifying the contributions of each amino acid in a given codon position.
@@ 2.2 Case study: Rapid punctuated evolution of SARS-CoV-2
PhASE TraCE was calibrated to the COVID-19 pandemic and the evolutionary trajectory of SARS-CoV-2 over the period of four years, from 2020 to 2023. The calibration explored ranges of over 90 parameters in four categories corresponding to modelling layers. Demographic parameters were census-calibrated [19, 20, 26, 25]. Most of the epidemiological and immunological parameters were calibrated in our prior studies [6, 24, 27, 25]. For example, parameter ranges related to non-pharmaceutical interventions were established based on prior studies [6, 24, 27, 25] and public health policies during the COVID-19 pandemic. In calibrating the natural history of the disease for this work, we used contemporary evidence, including studies of chronic infections [31, 30]). In contrast, all phylogenetic parameters were specifically calibrated for this study. The majority of phylogenetic parameters were determined using reported genomic evidence [18] and recent phylogenetic analyses [32, 17, 33, 34, 35, 36]. Some phylogenetic inputs that are specific to our model, such as the weight table quantifying the fitness contribution of amino acids, were calibrated by comparing the simulation outcomes corresponding to different parameterisations and selecting the most fitting parameterisations. The phylogenetic parameters are summarised in Table 4.
The results comprise simulations over a six-year period (2020 – 2025), across three artificial agent populations: 1.7 million agents (small), 8 million agents (medium), and 25.4 million agents (large), as described in S1 Text (Artificial agent-based population).
In calibrating PhASE TraCE, we examined COVID-19 pandemic and the size of susceptible, recovered and vaccinated populations (Capability 1); phylodynamic characteristics including transmissibility, fitness, the number of accumulated mutations relative to the ancestral strain, and genomic diversity (Capability 2); and the emergence of variants, analysed using the phylogenetic tree and statistical methods (Capability 3). The robustness of the model is established through sensitivity analysis, by varying key parameters and assessing their impact on simulation outcomes. See Table 4 for the list of phylogenetic parameters used in the SARS-CoV-2 case study and S1 Text (Sensitivity analysis) for more information on sensitivity analysis.
@@
Capability 1. Our results show recurrent incidence waves as illustrated in Fig 6A and 19, in concordance with empirical observations (Fig 1 and 19). Although the first simulated incidence peak (detected during 2021) is noticeably higher in larger populations, the incidence converged to around 2,000 cases per million after 2023, indicating endemicity in all simulated populations.
The main contributing factor to the recurrent incidence is the fluctuating number of recovered individuals (Fig 6B) that replenish the susceptible population after their immunity wanes. This subsequently leads to an increasing number of re-infection cases. Additionally, due to the diminishing immunity, declining vaccine uptake, and declining adoption of social distancing in the population (simulated to be decreasing from 50% of hosts in 2020 to 20% from 2024 onwards), the overall incidence slightly increases over the simulated timeframe after the first peak (Fig 6A). Notably, simulation results suggest that chronic infections strongly affect the magnitude of incidence oscillations (28). This finding may also explain the dampened oscillations observed in smaller populations. Emerging strains tend to dominate more strongly across a smaller population, resulting in: (i) reduced genomic diversity of the pathogen, (ii) a more homogeneous and consistent immune response across the population, and consequently (iii) a smoother incidence profile (see discussion of Capability 3 below).
We also observe an increasing number of individuals with multiple vaccination and infection records (Fig 6C). These dynamics slow down the transmission but are unable to completely eliminate the spread, due to the waning immunity and reduced vaccine efficacy against mutated variants.
Figs 6B and 6C only show dynamics for the 1.7 million population. We note that the pandemic patterns are consistent across different population sizes, as shown in Fig 20.
@@
Importantly, we observe a concurrent punctuated increase (i.e., ‘jump’) in both transmissibility, i.e., fitness $K$ (Fig 7A), and accumulated mutations $\widehat{D}$ (Fig 7B) at the start of 2021. This is a pattern which is also observed in empirical data illustrated in Fig 2A. The timing of the simulated ‘jump’ aligns well with the first incidence peak in 2021, illustrated in Fig 6A. This coincidence can be explained by the accumulation of fitness-increasing mutations in chronically infected hosts, due to a higher selective pressure on pathogens evolving in these hosts; consequently generating a highly transmissible strain (i.e., a variant of concern) (see S1 Text: Infected host categories). This is confirmed by a comparison with a counterfactual scenario without chronically infected hosts in which the jump is not observed, as shown in Fig 28 and analysed in S1 Text (Counterfactual modelling). We also traced the fitness contributions of different genomic regions, by computing the fitness per codon (Fig 13). Evidently, the spike and epitope regions, which contribute to both transmissibility and antigenicity, have the highest fitness $K$ per codon (Fig 24) and the highest number of accumulated mutations $\widehat{D}$ per codon (Fig 24). Notably, this high contribution is observed despite the short genome length of this region (45 codons), indicating that the fitness gain is a result of combining the increased transmissibility and the higher immune escape. In comparison, the spike and non-epitope region (55 codons) which contributes only to transmissibility has lower fitness $K$ and accumulated mutations $\widehat{D}$ per codon.
Similar jumps are observed at later pandemic stages in a majority of individual simulation runs (see Figs 7C, and 21 to 23), but their timing differs across realisations, and so the average profiles shown in Fig 7A and Fig 7B smooth the saltations. The very first saltation is mostly aligned across the individual realisations, and so it is quite prominent in both $K$ and $\widehat{D}$. Fig 7C shows that the fitness $K$ and accumulated mutations $\widehat{D}$ are strongly temporally aligned (with Pearson correlation 0.990.990.99 over 2,1912,1912 , 191 data points), suggesting that saltations in transmissibility are produced by accumulated mutations.
@@
To explain the punctuated increase in fitness, we traced changes in the distribution of fitness contributions across the simulated genome (Capability 2(ii)) by comparing the relative frequencies of amino acids in simulated ancestral genome (Fig 8 (I)) and the evolved distributions in circulating genomes after six simulation years (Fig 8 (II)). In simulated dynamics, we observed a clear shift towards amino acids with positive fitness contributions, in accordance with empirical observations (Figs 2C and 2D). This observation indicates that the viral mutations are subject to selective pressure transitioning to a higher point in the viral fitness landscape. Section 4 and S1 Text section A.3 provide a detailed description of the fitness contribution method.